OpenStudy (anonymous):

Factor x2 + 7x + 10. A. (x + 2)(x + 5) B. (x + 1)(x + 10) C. (x + 1)(x + 9) D. prime

OpenStudy (anonymous):

to factor qudratics (an expression that has x raised to the second power and not higher) find 2 numbers that when you multiply them = 10 but also when you add them = 7 try 5 and 2 5*2=10 5+2=7 (x+5)(x+2)

OpenStudy (anonymous):

thanks